The Singer Division of Patio Enterprises currently earns $2.6 million and has divisional assets of $20 million. The division manager is considering the acquisition of a new asset that will add to profit. The investment has a cost of $3,387,000 and will have a yearly cash flow of $843,000. The asset will be depreciated using the straight-line method over a six-year life and is expected to have no salvage value. Divisional performance is measured using ROI with beginning-of-year net book values in the denominator. The company’s cost of capital is 11 percent. Ignore taxes. The division manager learns that he has the option to lease the asset on a year-to-year lease for $742,000 per year. All depreciation and other tax benefits would accrue to the lessor.

Required:

a. What is the division's residual income before considering the project?

b. What is the division's residual income if the asset is purchased?

c. What is the division's residual income if the asset is leased?
Answer #1
a Residual income before considering the project
Current Earning $ 2,600,000
Less: Cost of capital $ -2,200,000
Residual income $      400,000
Cost of capital = 20000000*11%
b Residual income if the asset is purchased
Current Earning $ 2,600,000
New Earning $      843,000
Less: Cost of capital $ -2,572,570
Less: Depreciation $    -564,500
Residual income $      305,930
Cost of capital = (20000000+3387000)*11%
c Residual income if the asset is leased
Current Earning $ 2,600,000
New Earning $      843,000
Less: Lease Rent $    -742,000
Less: Cost of capital $ -2,200,000
Residual income $      501,000
Cost of capital = 20000000*11%